Adding 2 in the age of son makes his age quarter of his father's age. After 5 years, his father's age become 10 times less than his son's age.find their present ages. Please quickly give me answer, I have an exam tommorow.

Answer by stanbon(75887)

Adding 2 in the age of son makes his age quarter of his father's age. After 5 years, his father's age become 10 times less than his son's age.find their present ages.
-------
Equations:
s + 2 = (1/4)f
f+10 = (1/10)(s+10)
----------------------------
Modify:
f = 4s + 8
Substitute:
4s+8 + 10 = (1/10)s + 1
-------
3.9s = -17
----
Comment:: That would result in a negative answer.
Please check your posting of the problem.
Cheers,
Stan H.